Determine the processing sequence that would result using the Slack per operation rule.Given the information in the following table Remaining Number of Job Time (Days) Due Date Operations Remaining Processing 8 6
Answer #1

Slack per remaining operations (S/RO) is a priority sequencing rule that determines priority by dividing the slack by the number of operations that remain, including the one being scheduled. Slack is the difference between the time remaining until a job’s due date and the total shop time remaining, including that of the operation being scheduled.

S/RO = (Time remaining until the due date - Total processing time remaining)/ Number of operations remaining

S/RO (Job a)= (8-5)/2 = 1.50

S/RO (Job b)= (5-6)/4 = -0.25

S/RO (Job c)= (10-9)/4 = 0.25

S/RO (Job d)= (12-7)/3 = 1.66

S/RO (Job e)= (10-8)/2 = 1.00

Using slack per operation rule, we will sequence the jobs starting with lowest S/RO will yield following sequence-

b -> c -> e -> a -> d